A debt of $10 000 will be amortized by payments at the end of each quarter of a year for 10 years. Interest is at j4 = 10%. Determine the outstanding balance at the end of 6 years.

The Steins buy a house and take out a $255 000 mortgage. The mortgage is amortized over 25 years with monthly payments at j2 = 9%. After 3 and a half years, the Steins sell their house and the buyer wants to set up a new mortgage better tailored to his needs. The Steins find out that in addition to repaying the principal balance on their mortgage, they must pay a penalty equal to three months’ interest on the outstanding balance. What total amount must they repay?

A sinking fund is established with monthly deposits of $50 earning an interest rate j12 = 12%. If the balance immediately after the nth deposit equals $1161.96, what was the balance immediately after the (n-1)st deposit?

A. $1150.46

B. $1123.08

C. $1111.96

D. $1100.95

E. $1092.58
Answer #1

1.
=FV(10%/4,4*6,PMT(10%/4,4*10,-10000),-10000)=5200.62129898116

2.
=FV(((1+9%/2)^(2/12)-1)*12/12,12*3.5,PMT(((1+9%/2)^(2/12)-1)*12/12,12*25,-255000),-255000)*(1+((1+9%/2)^(2/12)-1)*12/12*3)=248924.002338781

3.
=(1161.96-50)/(1+12%/12)=1100.9504950495